# PRIMAL: Physically Reactive and Interactive Motor Model for Avatar Learning

### v1

@@ Abstract
To build a motor system of the interactive avatar, it is essential to develop a generative motion model drives the body to move through 3D space in a perpetual, realistic, controllable, and responsive manner. Although motion generation has been extensively studied, most methods do not support “embodied intelligence” due to their offline setting, slow speed, limited motion lengths, or unnatural movements. To overcome these limitations, we propose PRIMAL, an autoregressive diffusion model that is learned with a two-stage paradigm, inspired by recent advances in foundation models. In the pretraining stage, the model learns motion dynamics from a large number of sub-second motion segments, providing “motor primitives” from which more complex motions are built. In the adaptation phase, we employ a ControlNet-like adaptor to fine-tune the motor control for semantic action generation and spatial target reaching. Experiments show that physics effects emerge from our training. Given a single-frame initial state, our model not only generates unbounded, realistic, and controllable motion, but also enables the avatar to be responsive to induced impulses in real time. In addition, we can effectively and efficiently adapt our base model to few-shot personalized actions and the task of spatial control. Evaluations show that our proposed method outperforms state-of-the-art baselines. We leverage the model to create a real-time character animation system in Unreal Engine that is highly responsive and natural.  [^1] Code, models, and more results are available at:
@@ 1 Introduction
There has been rapid progress in 3D human motion generation conditioned on text or existing motion sequences [70, 84, 34, 32, 88, 39, 42, 44, 41, 68, 73]. Despite the realism of such approaches, they are not good models of how real humans behave, in that they are not interactive. True autonomous characters need embodied intelligence to respond in real-time to commands and environmental inputs such as external forces. Thus, our goal is to create an interactive 3D agent that responds to discrete commands, continuous control signals, and external impulses. To feel “alive”, such an agent must be continuously on-line and reactive. It must move on its own yet be able to respond to commands. We introduce such an agent using a two-stage approach, in which we first train a low-level motor system from motion capture (mocap) data such that it models the dynamics of human motion. This provides a platform upon which to build complex interactive behavior in a second stage. This two-stage approach provides a base “motor system” that captures the important dynamics of human motion, generalizes beyond the training data, and supports the training of more complex behaviors.
@@
At a high level, our approach, called PRIMAL (Physically Reactive and Interactive Motor Model for Avatar Learning), is inspired by models of biological motion that posit the presence of “motor primitives” that support more complex, responsive, behaviors [52]. In our approach, however, these “primitives” are not discrete, but rather are embodied in an autoregressive diffusion model that is trained to capture human motion dynamics over short time intervals. Without any physics simulation, we observe that this approach is able to model physically realistic movements, including ground contact, even though we never explicitly model contact. This provides a data-driven approach to avatar motor control that achieves physical realism without physics simulation.
@@
We make two key observations that ground our approach. First, existing autoregressive motion models usually predict a future motion conditioned on a past motion. Since a motion sequence implicitly contains semantic information, this ensures the predicted motion is semantically related to the past. However, it raises the risk that the model overfits to the semantics in the training data when the training samples are limited (as they are with existing mocap datasets). Second, human motion is governed by physics over short time intervals and by semantics over long ones. We use different approaches to model these. Physics plays a role in time intervals on the order of 0.5 sec, or 15 frames at 30fps.
@@
In PRIMAL, the first stage models this short-term motion: given an initial condition of the body, including its joint locations and velocities, we predict the motion of the next half second. This assumes that, over a short time horizon, that the motion of the future is independent of the past, conditioned on the given boundary condition. With this assumption, we have vast amounts of motion capture data with which to train the model; every segment of 15 frames is available for training. We demonstrate that training our generative model in this way captures the dynamics of human motion without doing any physics simulation. Note that we do not explicitly model foot-ground contact, yet the model learns to produce movements with physically plausible contact.
@@
Motivated by the representation power of diffusion [66, 26, 67], the first stage is implemented as an autoregressive diffusion model, which is trained on diverse half-second motion segments in a fully unsupervised manner. The trained model not only generates realistic and autonomous motions in the 3D space, but also makes it straightforward to control the avatar in real-time. Given this foundation, we can perturb the model with external inputs. For example, we simply reset any joint velocities to new values and the model will adapt its future motion appropriately. We can also drive the avatar to reach desired speeds and directions, using a classifier-based guidance (CBG) [14] approach, for which the CBG gradients are analytically derived.
@@
High-level and long-term behaviors can be composed as a time sequence of consecutive motion segments. This decouples the learning of dynamics and semantics, enabling us to learn them individually in a two-stage paradigm with a pretraining stage (motor control) and the adaptation stage (behavior generation). Specifically, we propose a generic adaptation approach based on ControlNet [83]. Here, we consider two types of adaptation. One focuses on spatial target reaching and the other on few-shot semantic action generation. To improve the effectiveness of the control signal, we follow [44] to add the control embedding to individual blocks of the transformer.
@@
We demonstrate the benefits of our interactive avatar with several applications. First, we create an interactive 3D environment in Unreal Engine in which the agent exhibits perpetual motion (see Fig. 1). The agent’s movement is controlled with standard game controllers to move in any direction at varied speeds. Unlike traditional game technology, however, our model is fully generative and all the motions are uniquely generated in real time based on the commands. We introduce external forces (“pulling” with a “magnet” or an impulsive force that “zaps” the character) that dynamically change the agent’s behavior (see Supplemental Video). The impulses induce physical forces to which the agent must respond. The agent reacts promptly to interventions in a natural way, without any explicit physics-based training, showing that the first stage learns a generic motor model.
Second, we demonstrate how our two-stage approach enables training a person/style-specific model using a tiny amount of motion capture data captured with a cell phone. This illustrates how pre-training a foundational movement stage provides a basic “motor system” upon which to build complex behaviors.
@@
In summary, PRIMAL produces naturalistic human motion common to methods trained on mocap data together with the physical plausibility and generalization of physics-based methods. To build a motor system for interactive avatars, we propose a new form of autoregressive model with a two-stage pretrain-and-adaptation learning paradigm. This first stage trains the model to capture motion dynamics, providing a reactive model that runs in real time. For the second stage, we propose an efficient and generic approach to adapt the base model to different avatar-related tasks. The interactive nature of PRIMAL effectively “breathes life” into an avatar, making it ideal for applications in gaming, AR/VR, controllable video generation, and interactive agents. We make our trained model, inference code, and the Unreal demo available for research purposes.
@@ 2 Related Work
Offline motion generation. Enabled by large mocap datasets [50, 20, 59], generative motion models have advanced rapidly, including methods focused on text-to-motion [70, 84, 32, 44, 31, 13], human-scene/object interactions [34, 88, 39, 40, 68, 80, 22, 21, 4, 69, 79], spatial control [45, 86, 35, 81, 15, 62, 78], dance generation [42, 75, 43], and more. Specifically, diffusion models [66, 67, 26] are a promising approach for motion generation, due to their representation power and controllability. For example, MDM [70] proposes a diffusion model with the transformer [77] network, which is trained with text-motion paired data [20] and is able to generate high-fidelity motions given text descriptions. OmniControl [81] leverages ControlNet [83] to create a new adaptor, enabling spatial control of text-to-motion. MotionLCM [13] proposes a consistency model for real-time text-to-motion, and implements autoregressive generation as a spatial control task. Most of these approaches are offline methods and have motion length limitations, making them inappropriate for interactive avatar control. In contrast to previous text-to-motion methods, out approach requires no paired data.
@@
More similar to our approach is OMG [44], which proposes a pretrain-then-finetune strategy for text-to-motion. Its pretraining phase trains foundation models with up to 1 billion parameters only with unlabeled mocap data, and the fine-tune phase leverages a ControlNet-like adaptor with a mixture-of-experts. In contrast to our approach, OMG is offline and their pre-training process focuses on generating motions up to 300 frames long, much longer than our short time horizon. Unlike OMG, our model formulation is autoregressive, which not only generates unbounded and controllable motion in real time, but also makes the avatar reactive to external perturbations.
@@
Physics-based motion generation. A key problem of the data-driven methods above is that they often lack physical plausibility; e.g. feet slide, float, or penetrate the floor, and impacts are not realistic. To address this, many methods have been developed that exploit physical simulation [46, 57, 62, 22, 78]. Hassan et al. [22] employ reinforcement learning (RL) and adversarial imitation learning to produce life-like and physically plausible human-scene interactions. PACER++ [78] generates physically plausible pedestrian locomotion in driving scenarios with fine-grained body part control. Simulation can also be efficiently combined with generative models. For example, Physdiff [82] projects the generated motion into a physically plausible motion space via physical simulation, which is applied within the reverse diffusion process. CLoSD [71] uses an efficient autoregressive diffusion model, DiP, for planning and applies simulation for tracking, generating unbounded physically plausible motions in 3D. An advantage of RL approaches is that they can generalize to adapt to external forces. However, they rely on approximate body models and the resulting motions are not yet as natural as those of data-drive methods. Additionally, these approaches tend to learn a narrow set of policies rather than a generic human motion control system.
@@
Autoregressive unbounded motion generation. Autogregressive models are trained to generate future frames conditioned on past frames. GAMMA [86] is a VAE [37] model that generates motion primitives recursively and leverages RL to train a policy for spatial target reaching. WANDR [15] is a conditional VAE that generates motions to reach a 3D target with the wrist joint. DIMOS [88] uses scene-aware policies for motion-primitive generation, enabling the synthesis of continuous human-scene interactions. Recently, diffusion models have been used for autoregressive motion modeling. For example, CAMDM [11] uses a diffusion model that treats motion style, spatial control, and motion history as separate tokens, producing unbounded motions for realtime character animation. Jiang et al. [33] propose a diffusion model that is conditioned on the past 2 frames, as well as text, scene, etc., to generate human-scene interactions. AMDM [64] uses an autoregrssive diffusion model with an MLP denoiser. In addition to generating random motions, policy networks are trained with RL for motion control in various tasks. The DiP model proposed in CLoSD [71] is conditioned on the motion history of 20 frames and text tokens, and generates a 40-frame motion in the future. DART [89] uses an online generation method that has an autoregressive VAE tokenizer and a latent autoregressive diffusion model, modeling motion primitives and performing high-fidelity text-to-motion in real time. Latent space optimization and RL policies are used for motion control on the pre-trained model.
@@
Compared to existing autoregressive diffusion models that generate $N$ future frames based on $M$ past frames, our model formulation is aligned with the initial value problem of multi-body dynamics, generating motions based on an initial state of a single frame that includes joint positions and velocities. With this form, our model learns human motion dynamics from unstructured and unlabeled mocap data directly, without the support of physics or annotations. Despite the lack of explicit physics in the model we find that physical effects emerge after training. In addition, neither motion tokenizers nor scheduled sampling [5] schemes are necessary in our method, leading to a simple and straightforward training workflow.
@@ 3.1 Preliminaries
Diffusion models are widely use for tasks like motion generation. In this work, we follow the DDPM formulation [26]. Provided a clean sample $\mathbf{X}^{0}\sim p(\mathbf{X}^{0})$, the forward diffusion process gradually adds Gaussian noise to the sample in $T$ steps, and produces a sequence of samples at different noise levels, i.e. $\{\mathbf{X}^{1},...,\mathbf{X}^{t},...,\mathbf{X}^{T}\}$. Given $q(\mathbf{X}^{t}|\mathbf{X}^{t-1})=\mathcal{N}\left(\sqrt{1-b_{t}}X^{t-1},b_{t}\mathbf{I}\right)$ with $\{b_{t}\in(0,1)\}_{t=1}^{T}$ being a set of noise variance scheduler variables, it has
@@
During test-time, the reverse diffusion process can be controlled via classifier-based guidance (CBG) [14] and classifier-free guidance (CFG) [25]. With a pre-trained diffusion model and objective functions to optimize during test-time, CBG computes the gradients of the objectives at each denoising step, and hence does not support realtime applications in general. On the other hand, CFG trains the conditional and unconditional diffusion models jointly, and guides the reverse diffusion process via weighted averaging, which is usually much more efficient than CBG. With $G(t,\mathbf{X}^{t},\emptyset)$ being the unconditional model, CFG is performed at each denoising step by
@@ 3.2 Motion Representation
Each motion segment spans 0.5 seconds in this work. This time duration corresponds to the average hang time of human jumps [^2] Hang time [6] describes how long someone can stay in the air under their own power and is a function of gravity and how high a person can jump. or nearly half of a walking cycle; this is a time frame in which human motion could be governed by physics. Following the canonical coordinate system proposed by [87, 86], the body-centric canonical coordinate is located at the pelvis of the first frame, with X,Y, and Z pointing left, up, and forward, respectively. The neural network is trained with canonicalized motion segments.
@@ 3.3 Neural Networks
Architectures. We formulate our autoregressive denoising network as $G(t,\mathbf{X}^{t},\bm{x}_{0})$ with transformers [77, 56], and investigate two architectures that differ in how the initial state is conditioned, as illustrated in Fig. 2. The first architecture leverages in-context conditioning, in which the diffusion time embedding and the initial state embedding are first added and then concatenated with the embedding of the noisy motion segment. We use SiLU [24, 16, 60] in the feed-forward layers. The second architecture follows the DiT network [56] with adaptive layer normalization [58].
Figure 2: Illustrations of our network architectures. The diffusion network is formulated as $\hat{\bm{X}}^{0}=G(t,{\bm{X}}^{t},\bm{x}_{0})$. The tea-green layers contain trainable parameters and the orange blocks are non-learnable operations. The pink squares denote the tokens at individual frames.
Training. Unlike existing autoregressive diffusion models, e.g. [89, 71, 11], that leverage annotated training data, our model focuses on modeling the motion dynamics and is trained only with unlabeled short mocap segments. Therefore, the data limitation issue is effectively overcome, since we can collect a huge amount of segments from AMASS [50]. In addition, our motion model is directly trained in the motion space, instead of the latent space as in [32, 89]. No tokenizers are trained in our method.
@@
in which the predicted motion state at the $i$-th frame is denoted by $\hat{\bm{x}}_{i}=(\hat{r}_{i},\hat{\varphi}_{i},\hat{\theta}_{i},\hat{J}_{i},\hat{\dot{J}}_{i})$. The ground truth body shape $\beta$ are used.

### v2

@@ Abstract
We formulate the motor system of an interactive avatar as a generative motion model that can drive the body to move through 3D space in a perpetual, realistic, controllable, and responsive manner. Although human motion generation has been extensively studied, many existing methods lack the responsiveness and realism of real human movements. Inspired by recent advances in foundation models, we propose PRIMAL, which is learned with a two-stage paradigm. In the pretraining stage, the model learns body movements from a large number of sub-second motion segments, providing a generative foundation from which more complex motions are built. This training is fully unsupervised without annotations. Given a single-frame initial state during inference, the pretrained model not only generates unbounded, realistic, and controllable motion, but also enables the avatar to be responsive to induced impulses in real time. In the adaptation phase, we employ a novel ControlNet-like adaptor to fine-tune the base model efficiently, adapting it to new tasks such as few-shot personalized action generation and spatial target reaching. Evaluations show that our proposed method outperforms state-of-the-art baselines. We leverage the model to create a real-time character animation system in Unreal Engine that feels highly responsive and natural.  [^1] Code, models, and more results are available at:
@@ 1 Introduction
There has been rapid progress in 3D human motion generation conditioned on text or existing motion sequences [70, 84, 34, 32, 88, 39, 42, 44, 41, 68, 73]. Despite the realism of such approaches, they are not good models of how real humans behave, in that they are not interactive. Autonomous characters need to respond in real-time to commands and environmental inputs such as external forces. To feel “alive”, such an agent must be continuously on-line and reactive. It must move on its own yet be able to respond to commands. We introduce such an agent using a two-stage approach, in which we first train a low-level motor system from motion capture (mocap) data such that it mimics the dynamics of human motion. This provides a platform upon which to build more complex interactive behaviors in a second stage. This two-stage approach provides a base “motor system” that implicitly models the dynamics of human motion, generalizes beyond the training data, and supports the training of more complex behaviors.
@@
Our approach, called PRIMAL (Physically Reactive and Interactive Motor model for Avatar Learning), models low-level motor control using an autoregressive diffusion model that is trained to capture human motion movement over short time intervals. Without any physics simulation, we observe that this approach is able to model physically realistic movements, including ground contact, even though we never explicitly model or label contact. This provides a data-driven approach to learning avatar motor control that appears physically realistic without physics simulation.
@@
We make two key observations that ground our approach. First, existing autoregressive motion models usually predict a future motion conditioned on a past motion. Since a motion sequence implicitly contains semantic information, this ensures the predicted motion is semantically related to the past. However, it raises the risk that the model overfits to the semantics in the training data when the training samples are limited (as they are with existing mocap datasets). This limits generalization, resulting in a model that does not make a good foundation for general motor control. Second, human motion is dominated by physics over short time intervals and by semantics over long ones. We use different approaches to model each of these. We observe that physics plays a critical role in time intervals on the order of 0.5 sec, or 15 frames at 30fps. On longer time horizons, higher level goals dominate.
@@
In PRIMAL, the first stage models this short-term motion: given an initial condition of the body, including its joint locations and velocities at a single time instant, we predict the motion of the next half second. This assumes that, over a short time horizon, the motion of the future is independent of the past. With this assumption, we have vast amounts of motion capture data with which to train the model; every segment of 15 frames is available for training. We demonstrate that training our generative model in this way mimics the dynamics of human motion without doing any physics simulation. Note that, while we do not explicitly model foot-ground contact, the model learns to produce movements with physically plausible contact.
@@
Motivated by the representation power of diffusion models [66, 26, 67], the first stage is implemented as a diffusion model, which is trained on diverse half-second motion segments in a fully unsupervised manner. The trained model not only generates realistic and autonomous motions in 3D space, but also makes it straightforward to control the avatar in real-time. For example, we simply reset any joint velocities to new values and the model will adapt its future motion appropriately. We can also drive the avatar to reach desired speeds and directions, using a classifier-based guidance (CBG) [14] approach, for which the CBG gradients are analytically derived.
@@
High-level and long-term behaviors can be composed as a time sequence of consecutive atomic actions. This decouples the learning of movements and semantics, enabling us to learn them individually in a two-stage paradigm with a pretraining stage (motor control) and the adaptation stage (behavior generation). Specifically, we propose a generic adaptation approach based on ControlNet [83] in the second stage. Here, we consider two types of adaptation. One focuses on spatial target reaching and the other on few-shot semantic action generation. To improve the effectiveness of the control signal, we follow [44] to add the control embedding to individual blocks of the transformer.
@@
We demonstrate the benefits of our interactive avatar with several applications. First, we create an interactive 3D environment in Unreal Engine, in which the agent exhibits perpetual motion (see Fig. 1). The agent’s movement is controlled with standard game controllers to move in any direction at varied speeds. Unlike traditional game technology, however, all the motions are uniquely generated in real time based on the commands. We introduce external forces (“pulling” with a “magnet” or an impulsive force that “zaps” the character) that dynamically change the agent’s behavior (see Supplemental Video). The agent reacts promptly to interventions in a natural way, without any explicit physics-based training, showing that the first stage learns a generic motor model. Second, we demonstrate how our approach enables training a person/style-specific model using a tiny amount of motion capture data captured with a cell phone. The results illustrate how pre-training a foundational movement stage provides a basic “motor system” upon which to build complex behaviors.
@@
In summary, PRIMAL produces naturalistic human motion common to methods trained on mocap data together with the physical plausibility and generalization of physics-based methods. To build a motor system for interactive avatars, we use an autoregressive diffusion model with a two-stage pretrain-and-adaptation learning paradigm. This first stage trains the model to capture human motion over short time scales, providing a reactive model that runs in real time. The second stage uses an efficient and generic approach to adapt the base model to different avatar-related tasks. The interactive nature of PRIMAL effectively “breathes life” into an avatar, making it appropriate for applications in gaming, AR/VR, controllable video generation, and interactive agents.
@@ 2 Related Work
Enabled by large mocap datasets [50, 20, 58], generative motion models have advanced rapidly, including methods focused on text-to-motion [70, 84, 32, 44, 31, 13], human-scene/object interactions [34, 88, 39, 40, 68, 80, 22, 21, 4, 69, 79], spatial control [45, 86, 35, 81, 15, 61, 78], dance generation [42, 75, 43], and more. Specifically, diffusion models [66, 67, 26] are a promising approach for motion generation, due to their representation power and controllability. For example, MDM [70] proposes a diffusion model with the transformer [77] network, which is trained with text-motion paired data [20] and is able to generate high-fidelity motions given text descriptions. OmniControl [81] leverages ControlNet [83] to create a new adaptor, enabling spatial control of text-to-motion. MotionLCM [13] proposes a consistency model for real-time text-to-motion, and implements autoregressive generation as a spatial control task. Most of these approaches are offline methods and have motion length limitations, making them inappropriate for interactive avatar control.
@@
More similar to our approach is OMG [44], which proposes a pretrain-then-finetune strategy for text-to-motion. Its pretraining phase trains foundation models with up to 1 billion parameters only with unlabeled mocap data, and the fine-tune phase leverages a ControlNet-like adaptor with a mixture-of-experts. In contrast to our approach, OMG is offline and their pretraining process focuses on generating motions up to 300 frames long, much longer than our short time horizon. Unlike OMG, our model formulation is autoregressive, which not only generates unbounded and controllable motion in real time, but also makes the avatar reactive to external perturbations.
@@
Physics-based motion generation. A key problem of the data-driven methods above is that they often lack physical plausibility; e.g. feet slide, float, or penetrate the floor, and impacts are not realistic. To address this, many methods have been developed that exploit physical simulation [46, 56, 61, 22, 78, 62]. Hassan et al. [22] employ reinforcement learning (RL) and adversarial imitation learning to produce life-like and physically plausible human-scene interactions. PACER++ [78] generates physically plausible pedestrian locomotion in driving scenarios with fine-grained body part control. Simulation can also be efficiently combined with generative models. For example, PhysDiff [82] projects the generated motion into a physically plausible motion space via physical simulation, which is applied within the reverse diffusion process. CLoSD [71] uses an efficient autoregressive diffusion model, DiP, for planning and applies simulation for tracking, generating unbounded physically plausible motions in 3D. InsActor [62] proposes an hierarchical approach to generate motions given instructions, which contains diffusion-based planning and simulation-based motion tracking. An advantage of simulation approaches is that they can generalize to adapt to external forces. However, the resulting motions are usually not yet as natural as those from diffusion models. Additionally, simulations tend to learn policies of a narrow set of motor skills, rather than a generic human motion control system.
@@
Autoregressive unbounded motion generation. Autogregressive models are trained to generate future frames conditioned on past frames. GAMMA [86] is a VAE [37] model that generates motion primitives recursively and leverages RL to train a policy for spatial target reaching. WANDR [15] is a conditional VAE that generates motions to reach a 3D target with the wrist joint. DIMOS [88] uses scene-aware policies for motion-primitive generation, enabling the synthesis of continuous human-scene interactions. Recently, diffusion models have been used for autoregressive motion modeling. For example, CAMDM [11] exploits a diffusion model that treats motion style, spatial control, and motion history as separate tokens, producing unbounded motions for real-time character animation. Jiang et al. [33] propose a diffusion model that is conditioned on the past 2 frames, as well as text, scene, etc., to generate human-scene interactions. AMDM [64] uses an autoregressive diffusion model with an MLP denoiser. In addition to generating random motions, policy networks are trained with RL for motion control in various tasks. The DiP model proposed in CLoSD [71] is conditioned on the motion history of 20 frames and text tokens, and generates a 40-frame motion in the future. DART [89] uses an online generation method that has an autoregressive VAE tokenizer and a latent autoregressive diffusion model, modeling motion primitives and performing high-fidelity text-to-motion in real time. Latent space optimization and RL policies are used for motion control on the pre-trained model.
@@
Compared to existing autoregressive diffusion models that generate $N$ future frames based on $M$ past frames, our model formulation generates motions based on an initial state of a single frame that includes joint positions and velocities. With this form, PRIMAL learns human movement from unstructured and unlabeled mocap data directly, without the support of physics or annotations. Despite the lack of explicit physics training, we find that visually plausible physical effects are present in the generated motions. In addition, neither motion tokenizers nor scheduled sampling [5] schemes are necessary in our method, leading to a simple and straightforward training workflow.
@@ 3.1 Preliminaries
Diffusion models are widely used for tasks like motion generation. In this work, we follow the DDPM formulation [26]. Provided a clean sample $\mathbf{X}^{0}\sim p(\mathbf{X}^{0})$, the forward diffusion process gradually adds Gaussian noise to the sample in $T$ steps, and produces a sequence of samples at different noise levels, i.e. $\{\mathbf{X}^{1},...,\mathbf{X}^{t},...,\mathbf{X}^{T}\}$. Given $q(\mathbf{X}^{t}|\mathbf{X}^{t-1})=\mathcal{N}\left(\sqrt{1-b_{t}}X^{t-1},b_{t}\mathbf{I}\right)$ with $\{b_{t}\in(0,1)\}_{t=1}^{T}$ being a set of noise variance scheduler variables, it has
@@
During test-time, the reverse diffusion process can be controlled via classifier-based guidance (CBG) [14] and classifier-free guidance (CFG) [25]. With a pre-trained diffusion model and the guidance objective functions, CBG computes the gradients of the objectives at each denoising step, and hence does not support real-time applications in general. On the other hand, CFG trains the conditional and unconditional diffusion models jointly, and guides the reverse diffusion process via weighted averaging, which is usually much more efficient than CBG. With $G(t,\mathbf{X}^{t},\emptyset)$ being the unconditional model, CFG is performed at each denoising step by
@@ 3.2 Motion Representation
Each motion segment spans 0.5 seconds in this work. This time duration corresponds to the average hang time of human jumps [^2] Hang time [6] describes how long someone can stay in the air under their own power and is a function of gravity and how high a person can jump. or nearly half of a walking cycle. This is a time frame during which human motion is dominated more by physical forces than by intentional control.
Following the canonical coordinate system proposed by [87, 86], the body-centric canonical coordinate is located at the pelvis of the first frame, with X,Y, and Z pointing left, up, and forward, respectively. The neural network is trained with canonicalized motion segments.
@@ 3.3 Neural Networks
Architecture.
Following e.g. [11, 64, 71, 89], in this paper the autoregressive diffusion model is defined as a diffusion model that generates $N$ future frames based on $M$ past frames. We formulate our denoising network as $G(t,\mathbf{X}^{t},\bm{x}_{0})$ with transformers [77, 55] and illustrate the architecture in Fig. 2. It leverages in-context conditioning, in which the diffusion time embedding and the initial state embedding are first added and then concatenated with the embedding of the noisy motion segment. SiLU [24, 16, 59] is used in the feed-forward layers.
Figure 2: Illustration of our network $\hat{\bm{X}}^{0}=G(t,{\bm{X}}^{t},\bm{x}_{0})$. The tea-green layers contain trainable parameters and the orange blocks are non-learnable operations. The pink squares denote the tokens at individual frames.
Training.
Unlike existing autoregressive diffusion models, e.g. [89, 71, 11], that leverage annotated training data, our model focuses on modeling the body movements and is trained only with unlabeled short mocap segments. Therefore, the data limitation issue is effectively overcome, since we can collect a huge amount of segments from AMASS [50]. In addition, our motion model is directly trained in the motion space, instead of the latent space as in [32, 89]. No separate tokenizers are trained in our method.
@@
in which the predicted motion state at the $i$-th frame is denoted by $\hat{\bm{x}}_{i}=(\hat{r}_{i},\hat{\varphi}_{i},\hat{\theta}_{i},\hat{J}_{i},\hat{\dot{J}}_{i})$. The ground truth body shape $\beta$ is used.